OpenStudy (kkutie7):

I Just have one question and a little bit of help. I'm going to write out the work I have and what the answer is supposed to be. I need to know where I went wrong.

OpenStudy (kkutie7):

A charge is fixed in place. A charge -q is released form rest a distance d from Q. Determine the velocity of -q as a function of its distance x from Q. You will need to write out Newton's 2nd law, apply the chain rule and then integrate both sides of the equation appropriately.

OpenStudy (kkutie7):

\[F = m \dfrac{dv}{dt}\] \[F=mv\frac{dv}{dx}\] \[\int Fdx=\int mvdv\] \[\int Fdx=\frac {mv^{2}}{2}\] \[\frac{Qq}{4\pi \epsilon_{0}}\int \frac{1}{x^{2}}dx=\frac{mv^{2}}{2}\] its obviously not x because the answer is this: \[v=\sqrt{\frac{Qq}{2\pi \epsilon _{0}m}(\frac{1}{x}-\frac{1}{d})}\]

OpenStudy (kkutie7):

@ganeshie8 think you could help me?